my father is 21 years older than me, after a time period of 12 years his age is twice of my age. how old am i ?
Answer by stanbon(75887)

my father is 21 years older than me, after a time period of 12 years his age is twice of my age. how old am i ?
---
Equations:
f = m + 21
(f+12)=2(m+12)
----
Substitute for "f" and solve for "m":
(m+21+12) = 2m+24
---
m = 9 yrs ("me's" age now)
==============================
Cheers,
Stan H.